### v1

# Training Gradient Boosted Decision Trees on Tabular Data Containing Label Noise for Classification Tasks

Abstract
Label noise refers to the phenomenon where instances in a data set are assigned to the wrong label. Label noise is harmful to classifier performance, increases model complexity and impairs feature selection. Addressing label noise is crucial, yet current research primarily focuses on image and text data using deep neural networks. This leaves a gap in the study of tabular data and gradient-boosted decision trees (GBDTs), the leading algorithm for tabular data. Different methods have already been developed which either try to filter label noise, model label noise while simultaneously training a classifier or use learning algorithms which remain effective even if label noise is present. This study aims to further investigate the effects of label noise on gradient-boosted decision trees and methods to mitigate those effects. Through comprehensive experiments and analysis, the implemented methods demonstrate state-of-the-art noise detection performance on the Adult dataset and achieve the highest classification precision and recall on the Adult and Breast Cancer datasets, respectively. In summary, this paper enhances the understanding of the impact of label noise on GBDTs and lays the groundwork for future research in noise detection and correction methods.
keywords:
Label noise , Gradient-boosted decision trees , Data quality , Data cleansing
journal: Information Processing & Management
[…]
[Debeka]organization=Debeka,city=Koblenz, country=Germany
[…]
[UK]organization=Universität Koblenz, Institute for Web Science & Technologies,city=Koblenz, country=Germany
[…]
[TUoS]organization=University of Sheffield,city=Sheffield, country=United Kingdom
[…]
[HSKo]organization=Hochschule Koblenz, Department of Math and Technology,city=Remagen, country=Germany
1 Introduction
In the area of machine learning and data-driven decision-making, accurate predictive modeling is crucial for diverse applications, from healthcare diagnostics to financial risk assessment. Among various techniques, gradient boosted decision trees (GBDTs) are the state-of-the-art algorithm for tabular data, consistently outperforming deep learning approaches (Grinsztajn et al., 2022).
However, real-world application of GBDTs poses several challenges. One particularly significant challenge is the presence of label noise in the data. Label noise refers to the phenomenon where instances in a dataset are assigned the wrong label. This differs from feature noise, which refers to inaccuracies in the observed feature values of an instance, rather than errors in the assigned labels (Zhu and Wu, 2004). While both feature and label noise can impair model performance, label noise is especially detrimental due to its strong impact on learning, given that there is only one label in contrast to multiple features (Frénay and Verleysen, 2014). Reliable labels are often costly and time intensive to acquire (Frénay and Verleysen, 2014). Although label noise detection is closely related to outlier and anomaly detection, it’s important to note that a mislabeled instance isn’t necessarily an outlier or anomaly and vice versa.
Label noise naturally occurs when human experts are involved (Frénay and Kaban, 2014). The provided information may be insufficient to perform reliable labeling, the expert may simply have made a mistake or the labeling task may leave room for interpretation, so that there is variability in the labeling between experts. Sometimes, communication or encoding problems can also be the cause of incorrect labels.
Feature noise as well as label noise deteriorates classification performance. Additionally, label noise can increase model complexity, increase the number of necessary training instances and complicate the identification of relevant features. Due to these problems, model comparison can be biased if the test set is polluted as well (Frénay and Verleysen, 2014).
1.1 Research Objective
Even though GBDTs outperform deep neural networks (DNNs) on tabular data (Grinsztajn et al., 2022), there are few studies exploring the effects of label noise on GBDTs and means of correcting for label noise (Ponti et al., 2022; Xiang et al., 2020; Brophy et al., 2023). Most of the new label noise research instead focuses on the application of DNNs on image or text data (Li et al., 2020; Jindal et al., 2017), despite the fact that tabular data is the most common data type (Shwartz-Ziv and Armon, 2022).
Therefore, the aim of this work is to
  • 1.
    provide an overview of state-of-the-art methods for detecting and handling label noise
  • 2.
    improve upon current state-of-the-art methods by implementing label-noise robust GBDT model
  • 3.
    evaluate the performance of the label-noise robust GBDT model by comparing it to state-of-the-art classifiers
1.2 Structure
To offer an overview of the state of label noise research and to set the stage for this work’s contribution, this section is followed by a review of related work (Section 2). Next, the “Preliminaries” section (Section 3) provides a formulation of the problem of training a supervised machine learning model on data containing label noise, the theoretical background for training GBDTs as well as a taxonomy of label noise. After that, the “Methodology” section (Section 4) outlines the implemented noise detection methods, how they are combined with noise correction methods and details the datasets and evaluation metrics used in this work. The “Experiments” section (Section 5) describes the experimental setup including how label noise is introduced into the data and which programming libraries and model hyperparameters are used. This section initially explores the impact of label noise on GBDTs and subsequently evaluates the performance of various noise detection and correction methods. The results from the experiments are then analyzed and discussed in the “Results and Discussion” section (Section 6). Finally, the “Conclusion” section (Section 7) synthesizes the findings, summarizes this work’s essential contributions and highlights avenues for future research.
2 Related work
This section presents research related to label noise and first covers the three main types of approaches when dealing with label noise: label noise robust models, label noise tolerant learning algorithms and data cleansing methods (Frénay and Verleysen, 2014). One paragraph is dedicated to the current landscape of label noise research, which involves training DNNs on text or image data. The strengths and weaknesses of each approach are also discussed. After that, the focus is drawn to research investigating training GBDTs on tabular data containing label noise as well as the fact that this type of research is limited.
The last section details how this work aims to fill that gap and derives the primary research questions of this work. It highlights the importance to understand the effects of label noise on GBDTs, to assess the applicability of deep neural network label noise detection methodologies on GBDTs, and to gauge the potential of these adapted GBDTs on benchmark datasets.
2.1 Label Noise Robust Models
Label noise robust models are based on algorithms that aren’t gravely affected by label noise in the first place. These models remain effective even if label noise isn’t cleaned or modeled. Dietterich (2000) showed that bagging generally performs better than boosting when label noise is present. Adaboost is a gradient boosting algorithm that increases the instance weights of misclassified instances (Géron, 2019).
The idea of label noise robust models promises that no or few considerations for label noise need to be made to train an effective model. Those few considerations usually come in the form of label noise robust loss functions and conventional regularization techniques. In practice, however, those models aren’t really that robust to label noise and turned out to only be adequate for simple noise settings and datasets. Also, these methods often have difficulty differentiating mislabeled from hard to learn training samples (Veit et al., 2017).
2.2 Label Noise Tolerant Learning Algorithms
Label noise tolerant learning algorithms either model label noise or regularize models to reduce their sensitivity to noisy instances. Often, the label noise model and classifier are trained simultaneously. Probabilistic approaches assume instances to be generated according to several probability distributions (Frénay and Verleysen, 2014). Bayesian probabilistic approaches use Bayesian priors to model the mislabeling probabilities (Gaba, 1993). But the priors should be chosen carefully, because the results depend on the quality of the chosen prior distribution. Frequentist approaches often use mixture models, where the data is assumed to originate from a mixture of a normal distribution and an “anomalous” distribution (Mansour and Parnas, 1998). Han et al. (2007) sort instances by their prediction scores and assume instances with the most mixed degree of positive and negative instances to be noisy and remove them using a revised EM-algorithm.
Probabilistic and frequentist approaches are grounded in a more theoretical approach than robust or data cleansing methods. But formulating the distribution of noisy and true labels based on specific assumptions may limit the effectiveness with more complex label noise (Li et al., 2019).
2.3 Data Cleansing Methods
Data cleansing methods aim to remove or relabel mislabeled instances to improve the performance of subsequent classifiers. One idea is to implement measures of anomaly (Sun et al., 2007) or measures of the effect on model complexity (Gamberger and Lavrac, 1997) and remove instances that exceed a certain threshold. Jeatrakul et al. (2010) simply remove instances misclassified by a classifier and Verbaeten and Assche (2003) discard instances with too large weights. There are also many heuristics based on k-nearest-neighbors or graphs. For example, removing instances whose label is different from the majority label in its k-neighborhood (Wilson, 1972).
Removing mislabeled instances rather than trying to relabel tends to be more efficient (Cuendet et al., 2007), except when the dataset is imbalanced, because underrepresented classes are more likely to be removed (Frénay and Verleysen, 2014).
Deep learning research often makes use of the so-called “small-loss-trick”, where small-loss samples from the training set are assumed to be clean and selected to update the network robustly (Han et al., 2018; Jiang et al., 2018; Li et al., 2020). Zheng et al. (2020) measure the ratio between a DNN classifier’s confidence on the noisy label and its own label prediction and flip the label to its prediction if the ratio exceeds a certain threshold. Additionally, a theoretical error bound is provided. Pleiss et al. (2020) measure the area under the margin to determine whether an instance is noisy or clean.
Data cleansing methods are often based on simple heuristics which are easy to implement. However, there is a trade-off between relabeling and filtering. Relabeling may introduce more errors, which accumulate in the process of training. Removing might delete minority classes first or instances which are hard to learn but informative, i.e. instances near a decision boundary (Johnson and Khoshgoftaar, 2022).
2.4 Deep Learning and GBDTs
In deep learning, common practices include adding robust architectures that model the noise transition matrix, regularization, improved loss functions and data cleansing, where instances with a likely true label are chosen to update the network (Song et al., 2020). To mitigate the accumulation of errors when further training is based on the network’s own predictions, some researchers have investigated the use of multiple networks (Chen et al., 2023). Co-teaching (Han et al., 2018) involves two neural networks, where each one selects low-loss instances for the other to use in backpropagation. Decouple (Malach and Shalev-Shwartz, 2017) updates the parameters of two networks only when their predictions are in disagreement to avoid overfitting to noisy instances. DivideMix (Li et al., 2020) operates by having each network discard the labels of high-loss instances and then provides the resulting partially-labeled data to the other network for semi-supervised training.
Deep learning achieves state-of-the-art performance on image and text data, but the methods often suffer from slow convergence, high computational cost due to an increased number of hyperparameters or large models due to deep architectures (Song et al., 2020).
Most modern label noise research is applied to image and text datasets using deep neural networks (DNNs) (Li et al., 2020; Jindal et al., 2017) and few on tabular datasets like those mentioned below. Architectures like TabNet (Arik and Pfister, 2021) have been developed to handle tabular data using deep neural networks, which outperform fully connected networks and other networks on several benchmarks, but still don’t achieve state-of-the-art performance as GBDTs do (Borisov et al., 2021).
So far, the research on applying GBDTs on data containing noisy labels is rather sparse, some exceptions include the following: Ponti et al. (2022) calculate the confidence, correctness and variability of the model’s predictions during the GBDT training process to detect and remove mislabeled instances, thereby improving the quality of the dataset. Brophy et al. (2023) adapt three existing methods for estimating the influence of training samples on predictions to GBDTs, which can be used to identify and correct mislabeled or problematic training samples. Xiang et al. (2020) use random label flipping and random feature subspaces to improve label noise robustness and increase the diversity of GBDTs and then aggregate them using bagging.
2.5 Summary and Research Questions
Common practices for dealing with label noise include modeling the noise, increasing the model’s robustness through regularization or robust loss functions or cleaning the data by removing, relabeling or reweighting noisy candidates. DNN-based techniques have received the most attention in recent years and therefore reached state-of-the-art performance on text and image datasets.
Tabular data is the most common data type in real-world applications (Shwartz-Ziv and Armon, 2022), but hasn’t been thoroughly studied in the context of learning with label noise. And tree-based methods like GBDTs remain the best algorithms for learning from tabular data (Grinsztajn et al., 2022). Therefore this paper studies the intersection of these two areas, examining effective strategies for learning from tabular data with label noise using GBDTs, with a focus on data cleansing methods and classification tasks.
Ponti et al.’s (2022) work focuses on data cleansing to combat label noise when training GBDTs. So far only the removal of noisy candidate instances was considered. Therefore this work continues their work by applying automatic relabeling of noisy candidate instances. Zheng et al.’s (2020) and Pleiss et al.’s (2020) data cleansing methods are widely cited and were selected to be adapted to GBDTs. The small-loss-trick (Li et al., 2020) and identifying noisy instances using the weights of boosting algorithms (Verbaeten and Assche, 2003) are also combined into a novel label noise detection method. All the above-mentioned approaches for noisy label detection and correction are implemented to improve the performance of GBDTs on data containing label noise.
Thus, this work aims to answer questions which are concerned with : (1) studying the effects of label noise on conventional GBDTs and (2) comparing the performance of different noise treatment methods implemented for this work. Regarding the first category, the questions are as follows:
  • RQ1. How many epochs of training does it take for noise to negatively impact model performance, i.e. how label noise robust are GBDTs naturally?
  • RQ2. What types of errors does the model make as training progresses?
  • RQ3. Are there certain types of noise which are harder to learn with?
For the second category, the research questions are as follows:
  • RQ4. How well do the detection methods detect noise?
  • RQ5. How do the detection methods change dependent on the noise rate?
  • RQ6. How do the detection metrics change over the course of training?
  • RQ7. Which correction method yields the highest model performance?
3 Preliminaries
This section provides an overview of the relevant theoretical foundations on which this paper builds upon. Each section focuses on a subject area which is needed to answer the research questions specified above.
The first section explains the basic idea of gradient boosting and how GBDT ensembles apply it by fitting shallow decision trees to the residual errors of their predecessor.
Lastly, the three types of label noise as characterized by Frénay and Verleysen (2014) are detailed, along with their implications for classification models. The types of noise injected into datasets within this work are also mentioned, along with the reasons for their selection. This section serves as a primer, equipping readers with the necessary theoretical knowledge to navigate the subsequent sections.
3.1 Training a GBDT Ensemble
The basic idea of Gradient Boosting is to train predictors sequentially, with each one aiming to fit the residual errors made by its predecessor (Géron, 2019). In other words, the sum of multiple weak learners $m_{t}$, which in the case of GBDTs are shallow decision trees, is supposed to approximate the target $y$ (Brophy et al., 2023; Si et al., 2017):
$y=\sum_{t=1}^{T}m_{t}(x),$ (1)
which can also be expressed as an additive model $f^{t}$ which we stop training once it reaches adequate performance:
$y$ $=f^{t}(x)=f^{t-1}(x)+m_{t}(x),$ (2)
$\Longleftrightarrow m_{t}(x)$ $=y-f^{t-1}(x).$ (3)
This means that $m_{t}$ fits the residual of the previous model $f^{t-1}$. To avoid overfitting to outliers, $m_{t}$ is usually weighted by a parameter called the learning rate $\eta$:
$f^{t}(x)=f^{t-1}(x)+\eta\cdot m_{t}(x).$ (4)
To generalize to any loss function, the negative gradient $g$ of the loss function $l$ with respect to $f^{t-1}$ is instead optimized, and is commonly referred to as the “pseudo-residual”. Using the gradient has the added benefit that it points to the direction of the steepest slope, rather than merely reflecting the magnitude of the difference, as the residual does:
$f^{t}(x)$ $=f^{t-1}(x)+\eta\cdot(-g_{t-1}(x,y)),$ (5)
$g_{t}(x,y)$ $=\frac{\partial l(f^{t}(x),y)}{\partial f^{t}(x)}.$ (6)
If, as an example, the mean squared error is the loss function of choice, then the next weak learner $m_{t}$ would be trained to fit:
$L_{MSE}$ $=\frac{1}{N}\sum_{i=1}^{N}(y_{i}-f^{t}(x_{i}))^{2},$ (7)
$g_{t}(x_{i},y_{i})$ $=\frac{\partial L_{MSE}}{\partial f^{t}(x_{i})}=-\frac{2}{N}(y_{i}-f^{t}(x_{i})),$ (8)
$f^{t}(x_{i})$ $=f^{t-1}(x_{i})+\eta\cdot m_{t}(x_{i}),$ (9)
$=f^{t-1}(x_{i})+\eta\cdot(-g_{t-1}(x_{i},y_{i})),$ (10)
$=f^{t-1}(x_{i})+\eta\cdot\left[-\left(-\frac{2}{N}(y_{i}-f^{t-1}(x_{i}))\right)\right],$ (11)
$=f^{t-1}(x)+\eta\cdot\frac{2}{N}(y_{i}-f^{t-1}(x_{i})),$ (12)
$\Longleftrightarrow m_{t}(x_{i})$ $=\frac{2}{N}(y_{i}-f^{t-1}(x_{i})).$ (13)
This shows that in the case of using the mean squared error as the loss function, the negative gradient is proportional to the residual of the previous weak learner $y_{i}-f^{t-1}(x_{i})$.
3.2 Taxonomy of Label Noise
Label noise can be categorized into three types: Noisy completely at random (NCAR), noisy at random (NAR) and noisy not at random (NNAR) (Frénay and Verleysen, 2014; Song et al., 2020). The corruption process is typically modeled with a noise transition matrix $S\in[0,1]^{c\times c}$, where $S_{ij}:=p(\tilde{y}=j|y=i)$ is the probability of the true label $i$ being flipped into a corrupted label $j$ and $c$ denotes the number of classes. This means that, starting from a noise rate of 50% and up, misclassification is just as or more likely than the true class.
NCAR noise is totally random and depends on neither instance features nor its true class. This type of noise is inherently symmetric, often termed “symmetric noise,” as it equally distributes the probability of the true label being flipped to any other label. With a noise rate $\tau\in[0,1]$, the noise transition matrix $(S_{ij})$ is defined as $S_{ii}=1-\tau$ and $S_{ij}=\frac{\tau}{c-1}$ for $i\neq j$.
NAR noise, also known as class-dependent noise, is independent of instance features but depends on its class, which allows modeling of asymmetric label noise. In that case, the true label is more likely to be mislabeled into a particular label than others, i.e. $S_{ii}=1-\tau$, but $S_{ij}\neq S_{ik}$ for some $k\neq i,j\neq i$. In a stricter case, called pair noise, the true label is flipped only to one certain label. Thus NAR noise can model the case where similar classes are easily confused with one another. For example, a “bird” is more likely to be confused with a “plane” than a “dog”.
NNAR noise, also known as instance- or feature-dependent noise, is deemed to be the most realistic, because it depends on both the instance’s features and its class. This could manifest as instances being more likely to be mislabeled because they are similar to instances of another class or certain regions of the feature space being more likely to be mislabeled. And some classes might be more likely misclassified as with NAR noise. Therefore the corruption probability is defined as $p(\tilde{y}=j|y_{i},x)$ where the data feature $x$ now also affects the chance of the instance being mislabeled.
As an example, Figure 1 shows how the noise transition matrices would look like on a dataset with four classes with no noise, 20% symmetric and 20% pair noise, respectively. In the first noise transition matrix, there is a 0% chance for each class to be switched to another class (excluding the true class). In the second noise transition matrix, there is about a 6.7% chance for each class to be switched to each other class, or a 20% chance in total to be switched to another class. In the last noise transition matrix, there is a 20% chance for each class to be switched to the next class, but a 0% chance to be switched to any other class.
Figure 1: Noise transition matrices for a dataset with four classes on no noise, 20% symmetric and 20% pair noise, respectively.
Given the challenges in modeling NNAR noise, as evidenced by the varied approaches to generating it (Zhang et al., 2021; Menon et al., 2018), and to increase the comparability to other works, this work focuses on introducing NCAR and NAR noise to the selected datasets. In the following sections, the noise types will be referred to as “symmetric” and “pair” noise, respectively.
4 Methodology
This section addresses the different steps taken to prepare the experiment settings with the goal of answering the research questions. First, four methods aimed at the detection of noisy instances are introduced and mathematically defined, followed by two label noise correction methods. Implementation details are also explained in these sections. After that, the datasets used for evaluation are listed and their characteristics briefly summarized, as well as the preprocessing measures taken to prepare the data for modeling. Lastly, the evaluation metrics used in this work and the reasoning behind that selection will be provided.
4.1 Implemented Methods
As mentioned in the “Related Works” section (Section 2), LRT-Correction and AUM Ranking were originally developed for DNNs and were adapted for this work to work on GBDTs. As opposed to Random Forests, which train fully grown trees on subsets of the data and aggregate their predictions using hard or soft voting, GBDTs fit shallow trees to the residual error of the current model, which is usually approximated using the gradient of the current model. Therefore, GBDTs optimize a loss function just as DNNs do, which is why DNN methods can be transferred to GBDTs.
Table 1 summarizes all implemented detection methods and some of their characteristics.
Method Original Domain Data
LRT-Correction DNN Past predicted class probabilities
Training Dynamics Statistics GBDT Class probabilities and predictions
AUM Ranking DNN Past raw margins
Gradients DNN Past gradients
Table 1: An overview of the implemented noise detection methods.
AUM-Ranking (Pleiss et al., 2020) and LRT-Correction (Zheng et al., 2020) use past predictions and calculate a metric related to the consistency of those predictions or a ratio between the predicted and the true class. Gradients isn’t based on a particular method, a combination of the small-loss-trick (Li et al., 2020) and the phenomenon that boosting algorithms tend to assign large weights to noisy instances (Verbaeten and Assche, 2003). These methods were chosen because they all apply to data cleansing, which is within the scope of this work as defined in Section 2.5, and are all well-known in the field of learning with label noise. Training Dynamics Statistics (Ponti et al., 2022) was adapted to an online algorithm, meaning that the metrics are computed at each training iteration instead of once at the end of training and then retraining a model based on these metrics. This change was made to investigate how the metrics change during training and for comparability to the other noise detection methods. Training Dynamics Statistics was chosen because it is one of the few works that investigate learning with label noise using GBDTs and focuses on data cleansing.
Though some methods additionally implement loss correction methods or similar measures, this work will solely focus on noise detection methods in combination with the relabeling and removal of suspicious instances. Also, despite the fact that some of the used datasets are imbalanced, taking class imbalance into account is beyond the scope of this work.
4.1.1 LRT-Correction
LRT-Correction stands for “likelihood ratio test correction” and is an algorithm that is based on a theorem provided by the authors quantifying how a classifiers prediction based on noisy data correlates to the purity of an instance’s label. Let $p(i|x)$ be the classifier’s predicted probability of class $i\in\{1,\dots,c\}$ and $\hat{y}=\mathrm{argmax}_{i}\ p(i|x)$ the predicted class. Based on the theorem, the ratio $LR(x,\tilde{y})$ between the classifier’s confidence on the noisy label $p(\tilde{y}|x)$ and its own label prediction $p(\hat{y}|x)$ is used to determine whether the instance should be relabeled (Zheng et al., 2020):
$LR(x,\tilde{y})$ $=\frac{p(\tilde{y}|x)}{p(\hat{y}|x)}$ (14)
If this ratio exceeds a certain threshold, i.e. the confidence on the predicted label $\hat{y}$ is much higher than that on the noisy label $\tilde{y}$, the label is flipped to the classifier’s prediction:
$\tilde{y}_{new}=\begin{cases}\hat{y},&\text{if }LR(x,\tilde{y})<\epsilon\\ \tilde{y},&\text{otherwise}\end{cases}.$ (15)
The network is trained using the cross-entropy loss $L_{CE}$. After a few warm-up iterations, an additional retroactive loss term $L_{retro}$ is added to enforce consistency of label predictions in consecutive epochs:
$l(f^{t}(x),\tilde{y},f^{t-1}(x))=L_{\text{retro}}(f^{t}(x),f^{t-1}(x))+L_{CE}(f^{t}(x),\tilde{y})$ (16)
$=\sum_{i=1}^{c}p^{t-1}(i|x)\log p(i|x)+\sum_{i=1}^{c}q_{\tilde{y}}(i)\log p(i|x),$
where $q_{\tilde{y}}(i)=[\tilde{y}=i]$ and where $p^{t-1}(i|x)$ is the probability of class $i$ as predicted by the previous model. The square bracket in $[\tilde{y}=i]$ is the Iverson bracket, which returns 1 if the condition is met and 0 otherwise. In this paper, the likelihood ratio $LR(x,\tilde{y})$ is used to identify noisy labels and either combined with LRT’s label flipping scheme as described in equation (15) or removing them. More on combining methods is described in Section 4.1.5.
4.1.2 Training Dynamics Statistics
Training dynamics describe the model’s behavior as its training progresses. In data cartography, statistics calculated based on training dynamics, i.e. the predicted probability $p(i|x)$ of class $i$ and predictions $\hat{y}$ across different estimators, are used to categorize data instances into “easy-to-learn”, “hard-to-learn” and “ambiguous”. Confidence $\mu(x)$ measures the average probability for the true label $\tilde{y}$ across all iterations. Variability $\sigma(x)$ measures the fluctuation of an instance’s true class probabilities during training and correctness $\gamma(x)$ defines the percentage of iterations in which the model correctly labels $x$ (Swayamdipta et al., 2020):
$\mu(x)$ $=\frac{1}{T}\sum^{T}_{t=1}p^{t}(\tilde{y}|x),$ (17)
$\sigma(x)$ $=\sqrt{\frac{\sum_{t=1}^{T}(p^{t}(\tilde{y}|x)-\mu(x))^{2}}{T}},$ (18)
$\gamma(x)$ $=\frac{1}{T}\sum^{T}_{t=1}[\hat{y}=\tilde{y}],$ (19)
with $T$ as the total number of trees (estimators).
Ponti et al. (2022) applied the idea of dataset cartography to GBDTs. They used the product of confidence and correctness to remove problematic instances either by threshold or an instance re-weighting scheme. In this work, we build upon the thresholding method, but in addition to removing problematic instances, also attempt to relabel them. The method will be denoted as ConfCorr in the Experiment section (Section 5).
4.1.3 AUM Ranking
“Area Under the Margin” (AUM) Ranking uses the margins at each epoch to differentiate easy, hard and mislabeled instances. The margin $M$ at epoch $t$ measures how much larger the assigned logit $z^{t}_{\tilde{y}}(x)$ is than the largest logit of another class and $\operatorname{AUM}$ is the mean margin of an instance over all epochs (Pleiss et al., 2020):
$M^{t}(x,\tilde{y})$ $=z^{t}_{\tilde{y}}(x)-\underset{i\neq\tilde{y}}{\max}\;z_{i}^{t}(x),$ (20)
$\operatorname{AUM}(x,\tilde{y})$ $=\frac{1}{T}\sum^{T}_{t=1}M^{t}(x,\tilde{y}).$ (21)
A positive margin indicates that a neural network is confident in its prediction. Conversely, a negative margin signifies a lack of confidence or even an incorrect prediction. In their work, Pleiss et al. (2020) additionally move threshold instances to an extra class so that the algorithm can learn to isolate mislabeled instances better from misclassified but correctly labeled instances. This work focuses on using the AUM metric as a means of identifying noisy instances.
4.1.4 Gradients
This method is based both on the “small-loss-trick” from deep learning as well as the growing weights of noisy instances in boosting algorithms (Verbaeten and Assche, 2003). With the small-loss-trick, small-loss samples from the training set are assumed to be clean and selected to update the network robustly (Han et al., 2018; Jiang et al., 2018; Li et al., 2020). Gradient boosting doesn’t assign weights to instances as standard boosting algorithms do, but the gradients with respect to each instance can be used instead.
An early experiment on the Covertype dataset confirms this: Figure 2 shows the maximum absolute gradient of clean and noisy instances per epoch. After just ten epochs the gradients of noisy instances are on average 1.5 times higher than the gradients of clean instances.
Figure 2: Maximum absolute gradients of noisy and clean instances per epoch (Covertype). The gradient is significantly larger for noisy instances.
In this work, the approach of Li et al. (2020) is followed, where they fit a two-component Gaussian Mixture Model (GMM) to the per-instance loss distribution over the last few epochs. The component with smaller mean is the clean probability. Since the selected objective function (see Section 5.1.2) returns the gradients with respect to each class per epoch, the largest absolute gradient of each instance was used to represent it.
4.1.5 Noise Correction
To limit the scope of this work, the correction methods were limited to removal and relabeling of noisy instances. As mentioned in the related works section (Section 2), other methods could include instance reweighting and loss correction.
All of the above methods were used for noisy instance detection and combined with either removal or relabeling for noisy instance correction, which amount to eight possible combinations. The percentage number of instances that may be removed was set to 80% so that the correction methods don’t accidentally remove almost all of the data, making further learning impossible. In the future, this percentage could be selected dynamically based on the estimated noise rate. For relabeling, the class with the highest probability averaged over the size of the history window (see Section 5.1.3) was chosen as the new label. Also, each instance could only be relabeled once to avoid error accumulation.
4.2 Data Sets
The model’s performance will be evaluated on four well-known public datasets, namely the UCI Covertype  [^1] https://archive.ics.uci.edu/dataset/31/covertype, accessed 30th July 2024, Dry Bean  [^2] https://archive.ics.uci.edu/dataset/602/dry+bean+dataset, accessed 30th July 2024, Adult  [^3] https://archive.ics.uci.edu/dataset/2/adult, accessed 30th July 2024 and Breast Cancer  [^4] https://archive.ics.uci.edu/dataset/17/breast+cancer+wisconsin+diagnostic, accessed 30th July 2024 dataset. They were chosen for comparability to other studies investigating the use of GBDTs on data containing label noise and because they are suitable for classification tasks.
The Covertype dataset contains cartographic descriptions of forest covers like elevation and soil type and is used to predict the type of forest cover (e.g. “pine”, “spruce” etc.). The labels were obtained from geospatial data from the US Forest Services. The features in the Dry Bean dataset were derived from high-resolution images of dry beans and are used to predict the type of the dry bean. The dry beans used for the images were obtained from certified seed producers (Koklu and Özkan, 2020), meaning that the data is professionally labeled. The Adult dataset is “a set of reasonably clean records” from the 1994 Census database and uses demographic information to predict whether a person’s income exceeds \$50K per year. Similar to the Dry Bean dataset, the Breast Cancer dataset contains features computed from images. The images display the fine needle aspirate (a procedure for obtaining cells from an internal organ or tumors) of breast mass and the features describe the characteristics of the cell nuclei present in the image. This information can be used to predict whether the samples contain benign or malignant cells, i.e. for cancer diagnosis.
Every dataset exhibits a certain level of class imbalance. For the Adult dataset, the class distribution is 76.07% to 23.93%. In the Breast Cancer dataset, the classes are divided into 37.26% and 62.74%. The Dry Bean dataset has its largest class representing 26.05% of the instances, while its smallest class accounts for 3.84%. Among these, the Covertype dataset is the most skewed, with the predominant class constituting 48.76% and the least common class comprising just 0.47% of the total data. The aforementioned proportions are visualized in Figure 3. The smallest class of the Covertype dataset consists of such few instances that its bar is barely visible in the figure.
Figure 3: The percentage of instances in the largest and smallest class per dataset. All datasets possess some degree of class imbalance.
The size of the datasets is also very diverse. The Covertype is the largest dataset, consisting of more than 500 thousand instances, whereas the Breast Cancer dataset merely contains 569. The specific dataset sizes can be viewed in Table 2 and are visualized in Figure 4.
Figure 4: The number of instances in each dataset. The size of the datasets varies a lot.
The labels of all public datasets are assumed to be almost clean, as they were either self-reported (e.g., the Adult dataset), based on other reliable datasets (e.g., the Covertype dataset), or labeled by professionals (e.g., the Dry Bean and Breast Cancer datasets). Characteristics of the datasets are summarized in Table 2.
Dataset #Instances #Features #Classes Types
Covertype 581012 54 7 Num
Dry Bean 13611 16 7 Num
Adult 48842 14 2 Cat/Num
Breast Cancer 569 30 2 Num
Table 2: Characteristics of the datasets used in this work.
All datasets were preprocessed by imputing missing data with the median of the feature if it was numeric and with mode otherwise. Numeric attributes were standardized to have a mean of zero and a standard deviation of one and categorical attributes were one-hot encoded. In some cases, features that leaked information about the target variable were discarded to prevent positively skewed results.
4.3 Metrics
In most papers, classification performance is evaluated using accuracy on the test set at various noise rates (Li et al., 2020; Xiang et al., 2020; Zheng et al., 2020). Here, accuracy on the train set is evaluated using the noisy labels and accuracy on the test set, since it isn’t polluted by noise, on the true labels only. Because this paper builds upon Ponti et al.’s (2022) work, precision, recall and the f1-score will also be included for greater comparability of the classification performances. Similarly, the noise detection performance will be evaluated using accuracy, the false negative rate (FNR) and the false positive rate (FPR).
5 Experiments
This section details the different experiment settings set up for answering the research questions and presents the information obtained in them. By providing clarity on the experimental procedures, this section ensures that the research is transparent and can be replicated by others. First, the experimental conditions are specified. This entails an explanation of the noise injection process, information about which noise types are used and the amount of noise that was injected into the data. The gradient-boosting library that was used as well as the model and other training parameters like early stopping are specified as well. The noise detection and correction methods used in the experiments are listed after that, which have already been introduced in Section 4.
The first section with experimental results covers the first set of research questions regarding the effects of label noise on GBDTs with no account for label noise. The experiments show that GBDTs are robust to label noise in the early epochs of training, especially to symmetric noise. They mostly predict the ground-truth true label even if it was altered due to label noise. Symmetric noise is generally harder to learn than pair noise but subsequent performance on the test set is higher.
The next section covers the second set of research questions, which are about the performance of the noise detection and correction methods and how they are influenced by the amount of noise present in the data. The experiments show that AUM and LRT perform best in terms of noise detection accuracy, closely followed by ConfCorr and Gradients. The percentage of instances marked as noisy by AUM and LRT can even be used to get a sense of the amount of noise present in the data. Combining these methods with either relabeling or removing instances marked as noisy results in a lower classification loss, but classification accuracy worsens. It could be that relabeling introduces more errors in the data and removing excludes hard to learn but informative instances from the training set.
To supplement the visual comparison of all applied methods, the “Final Results” section (Section 5.4) presents the results of different experimental settings in tabular form and compares them with other studies. AUM, LRT, and Gradients are demonstrated to achieve state-of-the-art noise detection accuracy on the Adult dataset. Additionally, ConfCorr and Gradients, when combined with noisy candidate removal, attain the highest recall score on the Breast Cancer dataset compared to other studies. ConfCorr, when combined with relabeling, achieved the highest precision score on the Adult dataset compared to other works. The most effective combination of noise detection and correction methods varies depending on the dataset, noise rate, and noise type. However, in terms of noise detection and classifier performance, the implemented methods were generally outperformed by other state-of-the-art work.
5.1 Experiment Conditions
5.1.1 Noise Injection
Assuming that all datasets are sufficiently clean, a noise transition matrix for pair and symmetric noise was applied to the training labels as defined in Section 3.2. The methods were evaluated on seven noise rates, $\tau\in\{0.0,0.1,0.2,0.3,0.4,0.5,0.6\}$, for each type of noise. In the case of pair noise, noise rates exceeding 50% make misclassification more likely than the correct label. However, these higher rates were still included to assess their impact on the model’s performance. In the final comparison (Section 5.4), only noise rates ranging from 10% to 40% were considered.
5.1.2 Model and Hyperparameters
For the classification task, a GBDT model was trained using the XGBoost library. The default parameters proved to be the most effective: maximum tree depth was set at 6, learning rate at 0.3, and the tree method was auto. Because the low-level boosting interface was used, early stopping had to be implemented manually with a minimum required loss improvement of 0.5 and a patience of 10 epochs. The model optimized the cross-entropy objective function and had a warm-up duration of 15 epochs. In deep learning, warm-up refers to the initial part of training where hyperparameters, especially the learning rate, are gradually increased from a lower value to the target value. This approach is employed to avoid instability or divergence of the training process, ensuring a more stable and effective convergence of the model. In the context of GBDTs, the term “warm-up” will be used to describe training the model for a few iterations to ensure that noise detection methods have enough past predictions to calculate their metrics from.
XGBoost’s softprob and logistic objective functions were used for multiclass and binary classification, respectively. softprob is essentially the cross-entropy loss function $H$, which is based on the softmax function $P_{\sigma}$ and returns the probability of each class per instance (Géron, 2019):
$P_{\sigma}(y=i)$ $=\frac{e^{z_{i}}}{\sum_{j=1}^{c}e^{z_{j}}},$ (22)
$H_{\sigma}(y)$ $=-\sum_{i=1}^{N}\sum_{k=1}^{c}[y_{i}=k]\cdot\log(P_{\sigma}(y_{i}=k)),$ (23)
where $z_{i}$ is the logit for class $i$. Logistic uses logistic regression as the objective. It is defined using the logistic function $P_{\kappa}$ to map the logits to a probability value between zero and one and again combined with cross-entropy $H_{\kappa}$ as the loss function:
$P_{\kappa}(y=1)=\frac{1}{1+e^{-z}},$ (24)
$H_{\kappa}(y)=\sum_{i=1}^{N}\left[y_{i}\log(P_{\kappa}(y_{i}=1))+(1-y_{i})\log(1-P_{\kappa}(y_{i}=1))\right].$ (25)
5.1.3 Algorithms
The detection methods LRT, ConfCorr, AUM and Gradients combined with the correction methods remove and relabel from Section 4 are compared with a baseline algorithm (denoted by none), which is a GBDT model trained without any noise treatment methods.
For both Gradients and ConfCorr, employing a two-component GMM to determine the optimal threshold at each iteration proved more effective in distinguishing clean from noisy instances than using a fixed or dynamic threshold. The threshold for LRT was based on its originating paper’s recommendations, setting the $\delta$ parameter to 1.0. Given that a positive margin indicates a correct prediction and a negative one an incorrect prediction, as per its original paper, the threshold for AUM was set to 0.
As in the original implementation, ConfCorr calculates its metrics incrementally across all epochs. The remaining noise detection methods use a history of five epochs to calculate their metrics from.
As outlined in Section 4.1.5, the removal of noisy labels was capped at 80% of the training data. If a detection algorithm identified more instances, the excess instances were retained. Each instance was allowed to be relabeled only once during relabeling to prevent the accumulation of errors.
5.2 Effect of Label Noise on Training GBDTs
This section deals with the first set of research questions, namely the effects of label noise on GBDTs without any methods correcting for label noise. To get a measure for the model’s classification performance, the performance of the trained model was evaluated on a clean test set, unpolluted by noise, and will be denoted as “test” in the following figures.
Both the Covertype and the Dry Bean dataset have more than two classes and were selected because it is generally easier to extrapolate findings from multiclass classification to binary classification than vice versa. The Covertype dataset is very large and complex, which is why it was chosen as the main dataset to generate the figures from. Using such datasets in research resembles real-world situations better, ensuring findings of this work aren’t just based on basic patterns but are more adaptable to varied scenarios.
However, sometimes the Covertype dataset was too large to calculate the metrics for all noise detection and treatment methods across all noise types and rates. In such cases, the Dry Bean dataset was used for these calculations instead. If there isn’t a significant difference in the plots generated for pair and symmetric noise, only one of them will be presented in this work.
To find out how naturally robust GBDTs are to label noise (RQ1) and what types of errors the model makes as training progresses (RQ2), the model was first trained without early stopping for 100 epochs and its predictions were collected and categorized into three groups: the prediction matches (1) the ground-truth label, (2) the noisy label or (3) another class altogether. Figure 5 shows that for both noise types, the model correctly predicts the true class most of the time at the start of training. The number of correct predictions decreases as training progresses and the number of predictions that match the noisy label grows. At about 60 epochs with 10% pair noise and 95 epochs with symmetric noise, predictions that match with the true and noisy label are equally likely. Only few predictions match a label that is neither the noisy nor the true label in the beginning and this number decreases as training progresses.
Furthermore, Figure 6 shows the classification accuracy on the train and test set per epoch. In both plots, the accuracy on the training set increases while that on the test set decreases as training progresses. When trained on pair noise, the decrease of accuracy on the test set and increase of the accuracy on the training set is much more rapid than when trained on symmetric noise. The increase in training accuracy is due to the model gradually adapting to the noisy data, as illustrated above. And because the model learns the noisy patterns, the test set accuracy decreases as training progresses. The test set accuracy is initially higher than the train set accuracy, because the test set accuracy is evaluated on the clean labels whereas the train set accuracy is evaluated on the noisy labels. The difference between the initial test and train set accuracy is approximately equal to the noise rate, i.e. 30%.
From those two figures one can infer that GBDTs, despite the fact that boosting algorithms are known to be very sensitive to noise, actually mostly predict the correct label and only gradually adapt to the noisy labels (RQ1, RQ2). Therefore early stopping is an important tool to use for training.
Figure 5: The types of predictions the model makes during training at 10% noise (Dry Bean). Only instances where the true label deviates from the noisy label are shown. The model predicts mostly the ground-truth true label in the beginning and gradually adapts to the noisy labels.
Figure 6: Classification accuracy on the train and test set per epoch at 30% noise without early stopping (Dry Bean). Test set accuracy on the clean test set decreases much slower when trained on symmetric noise, implying that GBDTs are more robust towards symmetric noise.
Figure 7 illustrates the model’s log loss per epoch, with each line representing training at a different noise rate. Lighter hues indicate lower noise rates, while darker hues represent higher ones. All training curves reach a minimum at approximately 8 epochs, with the darker lines converging at a significantly higher loss than the lighter ones. At 0% noise, the model achieves a similar log loss on both noise types. However, starting from noise rates as low as 10%, the final log loss when trained on pair noise is noticeably higher than when trained on symmetric noise. This indicates that model performance on the test set deteriorates as the noise rate increases and that pair noise presents a greater challenge for learning at noise rates higher than 10%. The reason behind this is that with pair noise, the label is flipped to only one class whereas with symmetric noise, the noise rate is equally distributed among the other classes. Therefore pair noise has a stronger influence when training the model.
Figure 8 plots the model’s accuracy against the noise rate it was trained on. For both symmetric and pair noise, the line has a downward slope, indicating that accuracy decreases as the noise rate increases. For pair noise, the line drops drastically at a noise rate of 50%, which is to be expected since from 50% pair noise upwards, predictions are random or misclassification more likely than the true class. However, prior to that point, the lines for symmetric and pair noise are quite similar, suggesting that training on either type of noise has a comparable impact on the model’s accuracy when early stopping is used.
Both Figure 7 and Figure 8 highlight the grave effect of label noise on classifier performance and the importance of investigating methods to counteract it.
An interesting point is that the slower decrease in test set accuracy when training on symmetric rather than pair noise (Figure 6) indicates that GBDTs are more robust to symmetric than pair noise.
Figure 7: Logloss on the test set per epoch at varying noise rates (Dry Bean). Logloss increases as the noise rate increases.
Figure 8: Classification accuracy on the test set at different noise rates (Dry Bean). Classification accuracy decreases as the noise rate increases.
Figure 9 shows the training curves when training on 30% symmetric and pair noise, respectively. With symmetric noise, performance on the training set converges earlier and at a higher logloss than with pair noise, indicating that symmetric noise is harder to learn with. On the other hand, final performance on the test set is in turn slightly higher with symmetric noise than with pair noise, signifying that the model would still perform well if the source of the label noise were to be removed or a label noise treatment method were to be employed.
Figure 9: Training curves at 30% noise (Covertype). Final test set performance is slightly higher when trained on symmetric noise.
Figure 10 visualizes the training curves on 10% and 40% pair noise, respectively. Train set performance converges at a much lower logloss with 10% noise than with 40% noise. Test set performance also converges at a lower logloss with 10% noise, but the difference in test set performance is smaller than in training set performance. The substantial difference in training set performance highlights the adverse impact of label noise on classification performance when evaluated on noisy data. Meanwhile, the smaller difference in test set performance suggests the robustness of GBDTs to label noise.
Figure 10: Training curves at 10% and 40% pair noise respectively (Dry Bean). The difference in performance is smaller on the test set than the train set, implying the model is also somewhat robust to pair noise.
In summary, while symmetric noise appears harder for the model to learn, it results in better logloss performance on a clean test set (RQ3). The training curves (Figure 6) confirm that, depending on when training is stopped, training on symmetric noise results in a higher classification accuracy on the test set than training on pair noise.
5.3 Performance Comparison
This section deals with the second set of research questions, which aim to compare the performance of the implemented noise detection and correction methods.
Figure 11 visualizes the metrics calculated by each of the implemented noise detection methods, separated by whether the instance the value was calculated for was noisy or clean. For effective differentiation between noisy and clean instances, it is desirable for the two distributions to be distinct and well-separated. The opposite would mean that the values calculated for noisy and clean instances don’t differ a lot and therefore the method couldn’t be reliably used to identify noisy instances.
The metric for AUM forms two close but distinct distributions of equal height, intersecting at a value of zero. The distributions for ConfCorr are widely separated; however, a significant number of clean instances fall within the noisy distribution, and a few noisy instances are found in the clean distribution. The Gradients distribution also forms two distinct distributions, but they are less peaked than those of AUM, with an intersection around 0.4. The clean distribution for LRT is tightly clustered around the value zero, although some noisy instances also fall within this range, with most distributed over values greater than zero.
From this, one might infer that AUM and Gradients could be the most effective in separating clean from noisy instances (RQ4), given their unimodal distributions. Their performance might be followed by ConfCorr, whose distributions are bimodal and have more overlap but still maintain a considerable gap. Although LRT clusters all clean instances closely around a single value, the presence of some noisy instances within this distribution suggests that LRT’s performance might be on the lower end.
Figure 11: Distribution of the calculated metrics by the ground truth noisiness of a label (Covertype), calculated on 30% pair noise. Gradients and AUM seem to separate between noisy and clean instances best.
To further investigate which methods detect noise best (RQ4), one can examine the confusion matrices comparing actual versus predicted noisy instances (Figure 12). All methods accurately identified noisy and clean instances approximately 80% of the time. Gradients achieved the highest performance, as it has the highest sum on the diagonal of the confusion matrix.
Figure 12: Label noise detection confusion matrices of each method on 30% pair noise (Covertype). Rows represent the ground truth and columns the predictions. Clean instances are represented by zero and noisy instances by one. All methods accurately identify noisy and clean instances approximately 80% of the time.
Figure 13 shows how the mean of the calculated metrics changes depending on the noise rate. AUM’s, ConfCorr’s and Gradients’ metric both converge as the noise rate increases, demonstrating that clean and noisy instances become more difficult to distinguish as the noise rate increases (RQ5). All metrics intersect and then diverge at a noise rate of 50% and beyond, which was expected since misclassification becomes as or more likely than the true class. The values calculated for clean instances remain close to zero with LRT and those for noisy instances approach zero as the noise rate increases. In the case of symmetric noise, the lines would intersect at a noise rate of $\frac{c-1}{c}$.
The convergence and eventual intersection of these lines demonstrate that for all methods, differentiating between clean and noisy instances becomes increasingly difficult as the noise rate approaches 50% (RQ5).
Figure 13: Mean metrics calculated at different noise rates with pair noise (Dry Bean). Most methods converge as the noise rate increases, meaning that clean and noisy instances become increasingly harder to differentiate.
Figure 14 plots the noise rate against noise detection accuracy. All methods achieve a noise detection accuracy of 70% or higher on both noise types. LRT and AUM overlap, which is why AUM’s line isn’t visible. They both also perform the best across all noise rates on symmetric noise, always achieving an accuracy of at least 90%. ConfCorr’s and Gradients’ accuracy is slightly worse for low and high noise rates. Pair noise degrades the quality of the data more quickly than symmetric noise, which is also reflected in the right plot of Figure 14. The correctness of the label is completely random at 50% pair noise, and likewise all methods reach a noise detection accuracy of about 50% and decreases further afterwards.
To summarize, LRT and AUM achieve the highest noise detection accuracy across all noise rates on the Dry Bean dataset, closely followed by Gradients and ConfCorr (RQ4).
Figure 14: Noise detection accuracy of all detection methods at various noise rates, calculated at the first epoch after warm-up (Dry Bean). All methods achieve a noise detection accuracy of 70% or higher on both noise types, with LRT and AUM achieving the highest accuracy of at least 90%.
Figure 15 maps the actual noise rate against the percentage of instances marked as noisy by AUM at the first epoch after warm-up. The black dotted line represents an ideal noise detection classifier, where the percentage of instances marked as noisy matches the actual noise rate present in the data. On the Dry Dean dataset, a few instances are marked as noisy even though no noise is present. For both pair and symmetric noise, the line stays very close to the “ideal” black dotted line. From a noise rate of 40% onward, AUM predicts about 35% of the instances to be noisy even as the noise rate increases. On the Covertype dataset, the two lines generally lie above the ideal, meaning that usually more instances are marked as noisy than are actually present in the dataset.
Therefore on simpler datasets with few outliers and little class imbalance, the percentage of instances marked as noisy may be used to approximate the amount of noise present in the data, especially if the noise is symmetric (RQ5).
Figure 15: Noise detection accuracy of AUM at various noise rates, calculated at the first epoch after warm-up. AUM marks about as many instances as noisy as are actually present in on the Dry Bean dataset, but tends to mark more as noisy as are actually present on the Covertype dataset. Therefore the proportion of instances marked as noisy can be used as an estimate for the noise rate on simpler datasets.
Similarly, Figure 16 shows the detection accuracy of all methods on the Dry Bean dataset at the first epoch after warm-up. LRT and AUM overlap again, which is why AUM’s line isn’t visible.
The most prominent difference to the description above is that Gradients’ percentage of marked instances drops rapidly at a pair noise level of 50%. Both Gradients and ConfCorr tend to label more instances as noisy than are actually present in the datasets. This could indicate that they are more sensitive to outliers or class imbalances. In contrast, LRT and AUM generally mark a number of instances as noisy that is close to the actual amount present in the data. Consequently, the percentage of instances marked as noisy by LRT and AUM can be used to approximate the level of noise present in simple dataset polluted by symmetric noise (RQ5). For pair noise exceeding 40%, these estimates become increasingly unreliable.
Figure 16: Noise detection accuracy of all detection methods at various noise rates, calculated at the first epoch after warm-up (Dry Bean). The percentage of instances marked as noisy by LRT and AUM is close to the actual noise rate.
During training, AUM, ConfCorr and Gradients all increase (Figure 17), but the distance between noisy and clean instances stays roughly the same. LRTs metrics decrease and converge as training progresses.
This means that metrics for noisy and clean instances tend to converge as training progresses (RQ6). Depending on the metric, both values either increase or decrease. The values calculated on noisy instances tend to have a steeper slope, meaning that they change more drastically during training than those calculated on clean instances.
The ability of all methods to discriminate between clean and noisy instances remains roughly constant during training, with the exception of LRT. The convergence for LRT’s plot shows that the model adapts to noisy instances as training progresses and therefore noisy and clean instances become increasingly harder to differentiate (RQ6), highlighting again that early stopping is an important tool to avoid overfitting noisy labels. For all methods, using a fixed threshold without early stopping will likely be ineffective in discriminating between clean and noise instances throughout the whole training procedure.
Figure 17: Metrics calculated at different epochs with 30% pair noise (Dry Bean). The ability to discriminate between clean and noisy instances remains roughly constant during training for most methods.
Figure 18 visualizes the training curves when removing or relabeling noisy candidate instances detected using the corresponding detection method. The black dotted line represents the baseline model, i.e. not correcting for label noise, and each detection method is represented by a colored line. For the duration of the warm-up period, which lasts 15 epochs, all methods perform the same. After that there is a significant drop in error for all detection methods, while the baseline model’s error stays high. But at the same time, accuracy decreases for all models.
With regard to the drop in error, it can be concluded that early stopping alone is not sufficient to robustly train a GBDT model on data containing noisy labels. Instead, removing or relabeling suspicious instances results in a significant drop in error. But the decrease in loss did not translate to an increase in accuracy. Though AUM combined with removal achieves a higher accuracy than the other methods, it’s still worse than the baseline model. In Section 5.4 we’ll look at the classification accuracy in diverse noise settings and different datasets to find out whether noise treatment actually worsens classification performance.
Removal and relabeling appear to perform roughly the same according to Figure 18. The drop in accuracy is slightly lower with relabeling as well as the increase in error after about twenty epochs. Based solely on this information, a definitive conclusion regarding the best correction method (RQ7) cannot be determined. A numeric instead of a visual comparison of all noise detection and correction methods in Section 5.4 will give a clearer perspective on how to definitively answer which noise correction method performs better in which situations.
Figure 18: Logloss and accuracy on the test set per epoch when removing or relabeling instances marked as noisy by a noise detection method, calculated on 30% pair noise (Covertype).
As an example, Figure 19 shows which instances were marked as noisy by AUM at each epoch. At the first epoch after warm-up, about 58% of the instances marked as noisy were in fact noisy. Since these instances are removed, resulting in a reduced training set size, only a few instances are marked as noisy in subsequent epochs, most of which are actually clean. When relabeling noisy candidates, a small number of instances are selected for relabeling each epoch, with about half of them being clean. However, even though no instances are removed during relabeling, each instance can only be relabeled once. This implies that, when using relabeling, more instances are identified as noisy, possibly due to the relabeled instances influencing the model’s perception of the remaining instances. On one hand, relabeling is beneficial as it utilizes the entire dataset, but on the other hand, it could introduce more errors and thus potentially degrade model performance. Regarding the instances marked as noisy by the other detection methods in the first epoch after warm-up, about 60% are actually noisy. The same goes for all the methods when trained on 30% symmetric noise (see online code).
As mentioned previously, a numeric comparison in the next subsection will clarify whether removing or relabeling suspicious instances reaches a better classification performance.
Figure 19: The types of instances that got marked as noisy by AUM per epoch, calculated at 30% pair noise (Covertype). About 40% of the instances marked as noisy are actually clean.
Instead of setting a fixed upper limit for the percentage of instances that can be removed, another approach is to define a threshold based on the top quantile of noise values, with instances above this threshold being marked for treatment. For example, when the threshold is set at the 90th percentile, the 10% of instances with the highest noise values are selected for treatment. Figure 20 shows the log-loss on the test set across different quantiles for various noise rates. The minima of the curves are relatively equidistant, suggesting that treating fewer instances than the actual proportion of noisy instances typically yields the best performance. This aligns with earlier findings showing that over one-third of the instances marked as noisy are, in fact, clean. Therefore, in cases where the level of noise can be estimated, adjusting the proportion of instances treated can lead to better outcomes. This method, however, is not employed in the results section and represents an avenue for further research.
Figure 20: Logloss per quantile of the top noise values to be treated. (Dry Bean). Treating fewer instances than the actual percentage of noisy instances tends to yield the best performance.
5.4 Final Results
A visual comparison between the implemented noise detection and correction methods can give a quick impression of their differences, but numeric data is needed for a more detailed and nuanced comparison. This section therefore presents the noise detection and classification performance results in various noise settings and on different datasets. Here, the performance of the implemented methods will also be compared to state-of-the-art research using the same metrics, datasets and similar noise settings. The following tables were generated at the most optimal epoch as indicated by early stopping. In them, the best values per category are marked in bold letters.
In their paper, Ponti et al. (2022) implemented a variant of pair noise. On the Adult dataset, they achieved a maximum noise detection accuracy of 86.93%, 83.97% and 82.58% at noise rates from 10% to 30% each. On the same dataset, the methods implemented in this work achieve a maximum noise detection accuracy of 99.70%, 99.20% and 98.56% at noise rates from 10% to 30% each (Table 3), thereby outperforming Ponti et al.’s work across all noise rates.
This work’s noise detection accuracy on the Covertype dataset on noise rates from 10% to 30% are 83.06%, 82.90% and 81.00% each (Table 4) and is thereby outperformed by Ponti et al’s work which reached a noise detection accuracy of 91.15%, 89.12% and 85.95%.
On the Breast Cancer dataset, Ponti et al. achieved a maximum noise detection accuracy of 95.57%, 93.68% and 89.68% at noise rates from 10% to 30% each. This work, though, underperforms on the Breast Cancer dataset, achieving a maximum noise detection accuracy of 92.06%, 90.75% and 85.25% at the three noise levels (Table 5).
The performance difference between this paper and Ponti et al.’s work when using ConfCorr may partially be due to the difference in the pair noise implementation mentioned above and in the way the threshold is chosen.
In this work, AUM and LRT performed best on the Adult dataset across all noise rates (Table 3), whereas on the Breast Cancer dataset, ConfCorr and Gradients are best at detecting noise (Table 5). On the Covertype dataset, ConfCorr reached the highest noise detection performance across all metrics. These findings demonstrate that the optimal noise detection method varies depending on the dataset it is applied to.
Noise Rate Detection Accuracy FNR FPR
0.10 AUM 99.70 0.29 0.29
ConfCorr 93.30 6.70 6.70
Gradients 94.00 6.02 6.02
LRT 99.70 0.29 0.29
0.20 AUM 99.20 0.80 0.80
ConfCorr 94.40 5.60 5.60
Gradients 98.00 2.03 2.03
LRT 99.20 0.80 0.80
0.30 AUM 98.50 1.48 1.48
ConfCorr 98.44 1.58 1.58
Gradients 98.56 1.46 1.46
LRT 98.50 1.48 1.48
Table 3: Noise detection performance on the Adult dataset with 10% to 30% pair noise. Highest values are marked by noise rate.
Noise Rate Detection Accuracy FNR FPR
0.10 AUM 80.94 19.10 19.10
ConfCorr 83.06 16.94 16.94
Gradients 77.20 22.80 22.80
LRT 80.94 19.10 19.10
0.20 AUM 80.90 19.16 19.16
ConfCorr 82.90 17.16 17.16
Gradients 79.90 20.11 20.11
LRT 80.90 19.16 19.16
0.30 AUM 79.60 20.34 20.34
ConfCorr 81.00 19.00 19.00
Gradients 78.80 21.16 21.16
LRT 79.60 20.34 20.34
Table 4: Noise detection performance on the Covertype dataset with 10% to 30% pair noise. Highest values are marked by noise rate.
Noise Rate Detection Accuracy FNR FPR
0.10 AUM 90.30 9.67 9.67
ConfCorr 92.06 7.91 7.91
Gradients 90.56 9.45 9.45
LRT 90.30 9.67 9.67
0.20 AUM 80.44 19.56 19.56
ConfCorr 90.75 9.23 9.23
Gradients 90.75 9.23 9.23
LRT 80.44 19.56 19.56
0.30 AUM 72.06 27.90 27.90
ConfCorr 85.25 14.73 14.73
Gradients 82.60 17.36 17.36
LRT 72.06 27.90 27.90
Table 5: Noise detection performance on the Breast Cancer dataset with 10% to 30% pair noise. Highest values are marked by noise rate.
It’s unclear whether the performance of the final classification results is calculated on a specific noise rate or averaged between the tested noise rates of 10% to 30% in Ponti et al.’s paper. If the former is true, it would skew their result in a more positive direction. On the Adult dataset, Ponti et al. achieved a maximum precision, recall and f1-score of 0.58, 0.89 and 0.70 each. Therefore this work achieved a higher precision (79.56%) than Ponti et al. but a lower recall (57.03%) and f1-score (66.06%)(Table 6).
On the Covertype dataset, Ponti et al. achieved a maximum precision, recall and f1-score of 0.78, 0.74 and 0.75 respectively. This work achieved a higher precision (78.80%) than Ponti et al. but a lower recall (62.44%) and f1-score (66.06%) (Table 7).
On the Breast Cancer dataset, Ponti et al. achieved a maximum precision, recall and f1-score of 1.00, 0.86 and 0.92 respectively. This work achieved a lower precision (94.10%) than Ponti et al. but a higher recall (90.25%) and about the same f1-score (91.56%) (Table 8) using ConfCorr or Gradients in combination with noisy candidate removal.
ConfCorr and Gradients performed best in terms of classification accuracy on the Adult dataset with 30% noise, especially when paired with relabeling as the noise correction method. On the Breast Cancer dataset, those methods performed best when combined with noisy candidate removal instead of relabeling. In the case of the Covertype datasets, attempting to correct for noise seemed to be counterproductive, potentially introducing more errors or eliminating informative instances. Not correcting for noise resulted in the highest classification performance in terms of accuracy, recall, and f1-score. The sole exception was ConfCorr combined with removal, which attained the highest precision score. This shows that the ideal noise correction method is dependent on both the dataset it is applied to and the noise detection method it is paired with.
Zheng et al. (2020) applied 10% pair noise in a way that doesn’t affect the class distribution, so the noise implementation again differs a little from this work. They tested their work on multiple small-sized datasets, from which one is used in this work as well. On the Breast Cancer dataset, they achieved a classification accuracy of 94.9%, outperforming this work by about 1.9% (Table 8).
Correction Detection Accuracy Precision Recall F1
none none 85.94 77.60 56.94 65.70
relabel AUM 86.00 78.40 56.22 65.50
ConfCorr 86.10 78.44 57.03 66.06
Gradients 86.10 78.44 57.03 66.00
LRT 86.00 78.40 56.22 65.50
remove AUM 86.00 78.44 56.20 65.50
ConfCorr 85.60 79.56 52.78 63.47
Gradients 85.50 78.75 52.70 63.16
LRT 86.00 78.40 56.16 65.44
Table 6: Classification performance on the Adult dataset with 30% pair noise.
Correction Detection Accuracy Precision Recall F1
none none 76.94 78.06 62.44 66.06
relabel AUM 73.75 76.56 56.16 59.44
ConfCorr 74.60 76.75 57.34 61.16
Gradients 74.50 76.60 57.16 60.72
LRT 73.70 76.30 56.20 59.44
remove AUM 74.44 78.00 58.70 62.90
ConfCorr 72.60 77.06 52.97 53.30
Gradients 73.30 78.80 52.12 53.06
LRT 74.44 78.00 58.60 62.75
Table 7: Classification performance on the Covertype dataset with 30% pair noise.
Noise Rate Noise Type Correction Detection Accuracy Precision Recall F1
0.10 pair none none 90.40 96.94 87.50 92.00
relabel AUM 91.25 94.30 91.70 92.94
ConfCorr 90.40 95.50 88.90 92.06
Gradients 90.40 95.50 88.90 92.06
LRT 90.40 95.50 88.90 92.06
remove AUM 91.25 94.30 91.70 92.94
ConfCorr 92.10 97.00 90.25 93.50
Gradients 93.00 97.06 91.70 94.30
LRT 91.25 95.56 90.25 92.90
0.30 none none 80.70 91.70 76.40 83.30
relabel AUM 79.80 90.20 76.40 82.70
ConfCorr 79.80 90.20 76.40 82.70
Gradients 79.80 90.20 76.40 82.70
LRT 79.80 90.20 76.40 82.70
remove AUM 78.94 91.40 73.60 81.56
ConfCorr 89.50 94.10 88.90 91.44
Gradients 89.50 92.90 90.25 91.56
LRT 78.94 91.40 73.60 81.56
Table 8: Classification performance on the Breast Cancer dataset with 10% and 30% pair noise. Highest values are marked by noise rate.
Table 9 displays the classification performance of all methods at different noise rates and types. On 20% pair noise, Gradients together with noisy candidate removal achieved the highest classification accuracy, precision, recall and f-1 score. On 20% symmetric noise, AUM combined with relabeling achieved the highest classification accuracy, but ConfCorr combined with removal the highest precision, recall (along with Gradients) and f1-score. On 40% pair noise, Gradients and noisy candidate removal reached the highest performance across all metrics (some on par with ConfCorr), but on symmetric noise, AUM combined with relabeling reached the highest classification performance metrics.
These results indicate that the most effective combination of noise detection and correction methods depends on the specific dataset, the type and the amount of noise present in the data.
Noise Rate Noise Type Correction Detection Accuracy Precision Recall F1
0.20 pair none none 90.60 91.40 90.56 90.94
relabel AUM 91.56 93.00 91.56 92.25
ConfCorr 91.70 93.06 91.80 92.40
Gradients 91.60 93.06 91.80 92.40
LRT 91.56 93.00 91.56 92.25
remove AUM 91.56 93.00 91.56 92.25
ConfCorr 92.00 93.60 93.06 93.30
Gradients 92.10 93.75 93.20 93.50
LRT 91.56 93.00 91.56 92.25
symmetric none none 91.50 93.00 92.10 92.56
relabel AUM 91.90 93.44 92.75 93.06
ConfCorr 91.75 93.44 92.70 93.00
Gradients 91.75 93.44 92.70 93.00
LRT 91.75 93.40 92.56 92.94
remove AUM 91.75 93.30 92.60 92.94
ConfCorr 91.75 93.50 92.90 93.20
Gradients 91.75 93.40 92.90 93.10
LRT 91.75 93.30 92.60 92.94
0.40 pair none none 78.25 75.50 76.56 75.94
relabel AUM 86.94 86.80 85.70 86.10
ConfCorr 86.90 86.44 85.70 86.00
Gradients 86.70 86.00 85.40 85.60
LRT 87.00 86.75 85.75 86.10
remove AUM 87.30 87.06 86.00 86.44
ConfCorr 88.30 88.44 87.00 87.56
Gradients 88.60 88.44 87.10 87.70
LRT 87.25 87.06 86.06 86.40
symmetric none none 91.00 92.25 91.56 91.90
relabel AUM 91.90 93.44 92.75 93.06
ConfCorr 91.56 93.06 92.40 92.70
Gradients 91.56 93.06 92.30 92.70
LRT 91.70 93.25 92.56 92.90
remove AUM 91.90 93.25 92.56 92.90
ConfCorr 91.20 92.56 92.00 92.30
Gradients 91.25 92.70 92.10 92.40
LRT 91.75 93.25 92.44 92.80
Table 9: Classification performance on the Dry Bean dataset with noise rates 20% and 40%. Highest values are marked by noise type.
6 Results and Discussion
This section summarizes the experimental findings, their implications, and evaluates the validity of the results in relation to the research questions posed in Section 2.5. The experiments showed that GBDTs are robust to label noise, especially symmetric noise, reaching a local minimum on the clean test set quickly, suggesting that early stopping can yield a good model.
The methods AUM and LRT achieved a noise detection accuracy of 99.30% on the Adult dataset. Differentiating noisy and clean instances becomes harder as noise increases. The metrics for clean and noisy instances change during training, so a dynamic threshold or early stopping is advisable. The effectiveness of noise detection and correction methods varies by dataset and noise type.
The Covertype dataset, with its multiple classes and high class imbalance, performed differently than the binary Adult dataset. Although LRT struggled with separating clean from noisy instances, it consistently assigned low values to clean instances, indicating effectiveness. AUM achieved the highest classification accuracy on symmetric noise on the Dry Bean dataset when paired with relabeling.
ConfCorr effectively separates clean from noisy instances, though its optimal threshold depends on noise rate, type, and dataset. A two-component GMM can work around this but is less efficient. ConfCorr and Gradients attained high classification precision on the Adult dataset and recall on the Breast Cancer dataset.
Though Gradients initially seemed less promising for noise detection, it achieved high precision and recall on different datasets. Similar to ConfCorr, it faces challenges with determining an optimal threshold.
In summary, the choice of detection and correction methods is influenced by noise rate, type, and dataset. Class imbalance may render some methods unsuitable, as they might remove instances from smaller classes. Relabeling can sometimes introduce more errors, while removing suspicious instances might discard informative data. It’s generally advisable to treat fewer instances than the estimated amount of noise present in the data. Multiple methods should be tested to find the optimal solution for label noise problems.
Due to time and computational constraints, each experiment was performed only once. The variability in method performance across different datasets suggests that favorable outcomes might be due to chance rather than inherent effectiveness. Cross-validation could address these concerns. Additionally, pair and symmetric noise may not represent real-world label noise, limiting the applicability of these findings.
7 Conclusion
This work explores the effects of label noise on gradient boosted decision trees (GBDTs) and compares different methods of label noise detection and correction. The primary objectives, as outlined in Section 1.1, are to provide an overview of the label noise research landscape, implement a GBDT which improves upon other GBDTs trained on noisy, tabular data and evaluate its performance on benchmark datasets.
The literature review provided a structured overview of label noise research, categorizing it by their general approach. Then it highlighted the most recent advancements in the field, namely the training of deep neural networks (DNNs) on noisy image and text data. Training GBDTs on tabular data was then identified as a notable gap in research. Stemming from this observation, the primary research questions were formulated. These encompass studying the impact of label noise on GBDTs, adapting certain methodologies from DNN research to GBDTs, and subsequently evaluating their effectiveness on benchmark datasets, with a comparative analysis against other state-of-the-art research.
This was achieved by first gathering data, preprocessing it, training GBDTs at various noise rates and types and finally interpreting the results both visually and numerically. Four benchmark datasets, namely the Breast Cancer, Covertype, Adult, and Dry Bean dataset, were used to evaluate the implemented methods. The first three datasets served as a basis for comparison with state-of-the-art work, while the Dry Bean and Covertype datasets helped explore the behavior of GBDTs under label noise conditions.
The experiments revealed that GBDTs are robust to label noise in the early stages of training, especially to symmetric noise, and can achieve good performance by utilizing early stopping. In the scope of this work, methods from deep learning were adapted and both AUM and LRT achieved state-of-the-art noise detection performance on the Adult dataset. The percentage of instances marked as noisy by AUM and LRT can be utilized to approximate the level of noise present in the data, provided that the dataset is not excessively imbalanced or contains too many outliers. But it is still advisable to treat fewer instances than the estimated amount of noise present in the data.
Inconsistency in performance of the methods across various datasets implies that the success of the implemented label noise detection and correction methods could depend on factors such as the dataset, noise rate, noise type, and the chosen correction technique. Alternatively, these favorable results might be attributed to random chance rather than the inherent effectiveness of the methods. Conducting repeated experiments and employing cross-validation could resolve these ambiguities.
Future research could explore other relabeling methods, beyond using the most confident or frequent predictions from recent epochs. And because high noise detection performance didn’t always translate in a high increase in classification accuracy, further measures like error correction may be needed to effectively make use of the noise detection methods. Since classification performance decreased when removing or relabeling noisy candidates on imbalanced datasets, future research could also investigate methods to reliably detect noisy instances on data with high class imbalance using GBDTs. Until now, GBDTs generally outperform DNNs on tabular data, but if that were to change it would be interesting to investigate whether label noise treatment methods from image and text classification are effective on tabular data as well.
In summary, this work conducted a comprehensive analysis of the impact of label noise on GBDTs, offering practical guidance. Two methods for dealing with label noise from the area of deep learning were successfully adapted to work with GBDTs. Building on the work of Ponti et al. (2022), their noise detection algorithm was combined with automatic relabeling of noisy candidate instances. These modified GBDTs were tested against benchmark datasets, and their performance was compared to current state-of-the-art approaches. The findings here provide a basis for continued exploration and improvement in handling label noise, particularly in the context of tabular data.
8 CRediT authorship contribution statement
Anita Eisenbürger: Conceptualization, Data Curation, Formal analysis, Investigation, Methodology, Visualization, Writing – original draft, Writing – review & editing. Daniel Otten: Supervision, Writing – review & editing. Frank Hopfgartner: Supervision, Writing – review & editing. Anselm Hudde: Writing – review & editing.